iosangel:

Can somebody walk me through this question I've watched a couple videos but still don't understand.. 4(3x^2 y^4)^3/(2x^3y^5)^4

iosangel:

https://prnt.sc/pnu56m

iosangel:

1 attachment
mhchen:

\[ \frac{ 4(3x^{2}y^{4})^{3} }{ (2x^{3}y^{5})^{4} } \] \[ \frac{ 2^{2}(3^{3}x^{6}y^{12}) }{ 2^{4}x^{12}y^{20} } \] \[ \frac{ 3^{3} }{ 2^{2}x^{6}y^{8} } \] Do you see what I did in each step?

akrouse6:

Step 1 : Equation at the end of step 1 : (((3•(x2))•(y4))3) 4•—————————————————— ((2x3•(y5))4) Step 2 : Equation at the end of step 2 : ((3x2 • (y4))3) 4 • ——————————————— 24x12y20 Step 3 : 33x6y12 Simplify ———————— 24x12y20 Dividing exponential expressions : 3.1 x6 divided by x12 = x(6 - 12) = x(-6) = 1/x6 Dividing exponential expressions : 3.2 y12 divided by y20 = y(12 - 20) = y(-8) = 1/y8 Equation at the end of step 3 : 27 4 • —————— 16x6y8 Step 4 : Final result : 27 ————— 4x6y8

Skatercat:

\frac{27}{4x^6y^8}

werewolf:

complicated